OpenStudy (laweejay):

Help!(Optimization)... Two towns A&B are 13.0 miles apart and located 8.0 and 3.0 miles east of a long, straight highway. A construction company has a contract to build a road from town A to the highway and then to town B. Determine the length of the shortest road that meets these requirements.

OpenStudy (dumbcow):

 the length of road = x+y using pythagorean thm \[x+y = \sqrt{a^2 +64} + \sqrt{(12-a)^2 +9}\] to minimize x+y, set derivative equal to 0 \[\frac{a}{\sqrt{a^2 +64}} + \frac{a-12}{\sqrt{(12-a)^2 +9}} = 0\] solving for a leads to a 4th degree polynomial \[a^2 (a^2 -24a +153) = (12-a)^2 (a^2 +64)\] find zero with graphing software \[a = \frac{96}{11}\] plug in to get min value for "x+y" \[\min_{x+y} = \sqrt{265}\]